OpenStudy (anonymous):

(square root of 2)^(x+4)=4^x I just want to know how to get the bases to be the same. Do you square both sides, or do I "arrange" the right side of the equation to get the bases to be the same?

OpenStudy (amistre64):

sqrt(2)^(x+4) = 4^x square both sides to get rid of the sqrt sign

OpenStudy (amistre64):

[sqrt(2)^(x+4)]^2 = [sqrt(2)^2]^(x+4) [4^x]^2 = [4^2]^x

OpenStudy (amistre64):

we get: 2^(x+4) = 4^2x

myininaya (myininaya):

I get x=4/3

myininaya (myininaya):

I wrote both sides as base sqrt(2)

myininaya (myininaya):

4=(sqrt(2))^4

myininaya (myininaya):

(sqrt(2))^(x+4)=(sqrt(2))^(4x)

myininaya (myininaya):

so x+4=4x